Lulabalula:

Washington Middle School sold 405 tickets to the annual musical production for Friday night. If the tickets sold for Friday night were only 45 percent of the total tickets sold, how many total tickets were sold for the production?

XioGonz:

Percent of tickets sold on Friday = 45% so 45 The Total tickets = 900 The Tickets sold on Friday = 45/100×900 = (45×900)/100 = 40500÷100 or 40500/100

XioGonz:

So what is 40500/100

TWALKER16:

900 let the total tickets sold be x, 45% of x=405, x=(100*405)/45, x=900

Lulabalula:

4050

Lulabalula:

4050

TWALKER16:

The answer would be 900

TWALKER16:

i gave an explantation

TWALKER16:

Can we please close this if we are done with it

Imagine:

@sammixboo can you please close the question?

Imagine:

Thanks!